# Integrating Large Language Models into Recommendation via Mutual Augmentation and Adaptive Aggregation

### v1

@@ 5. Experiment
In this section, we present a thorough empirical evaluation to validate the effectiveness of our proposed framework. Specifically, our objective is to investigate whether the incorporation of our proposed Llama4Rec could enhance existing recommendation models. The overarching goal is to answer the following research questions:
@@
  • RQ3: How do different hyper-parameters impact the overall performance of the framework?
5.1. Experiment Setup
5.1.1. Dataset
Following (Bao et al., 2023), we rigorously evaluate the performance of our proposed framework by employing three heterogeneous, real-world datasets. MovieLens [^1] https://grouplens.org/datasets/movielens/ (Harper and Konstan, 2015) serve as benchmark datasets in the realm of movie recommendation methods. We employ two variants of the dataset: MovieLens-100K (ML-100K) and MovieLens-1M (ML-1M). The former consists of approximately 100,000 user-item ratings, while the latter scales up to roughly 1,000,000 ratings. BookCrossing [^2] Due to the absence of timestamp data, we synthesize historical interactions through random sampling. (Ziegler et al., 2005) includes user-generated book ratings on a scale of 1 to 10, alongside metadata such as ‘Book-Author’ and ‘Book-Title’. We employ LLM to augment the ‘director’ and ‘star’ features for ML-100K and ML-1M datasets, and augment the ‘genre’ and ‘page length’ features for the BookCrossing dataset. To ensure the data quality, we adopt the 5-core setting, i.e., we filter unpopular users and items with fewer than five interactions for the BookCrossing dataset. The key characteristics of these datasets are delineated in Table 1.
Table 1. Dataset Description.
@@ 5.1.1. Dataset
5.1.2. Evaluation Metrics.
Aligning with (He et al., 2020; Sun et al., 2019), for the top-$k$ recommendation task, we turn to two well-established metrics: Hit Ratio (HR) and Normalized Discounted Cumulative Gain (NDCG), denoted by H and N, respectively. In our experiments, $k$ is configured to be either 3 or 5 for a comprehensive evaluation, similar to the experiment setting in (Zhang et al., 2023c). In accordance with (Fan et al., 2019), we employ Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) as evaluation metrics to ascertain the performance of the rating prediction task.
Table 2. Performance achieved by different direct recommendation methods.
Table 3. Performance achieved by different sequential recommendation methods.
@@
5.1.3. Data Preprocessing.
Following the methodology of prior works (Zhang et al., 2023c; Luo et al., 2022b), we adopt a leave-one-out evaluation strategy. More specifically, within each user’s interaction sequence, we choose the most recent item as the test instance. The item immediately preceding this serves as the validation instance, while all remaining interactions are used to constitute the training set. Moreover, regarding the instruction-tuning dataset construction, we randomly sampled 5K instructions for each recommendation task on the ML-100K, ML-1M, and BookCrossing datasets, respectively. We eliminated instructions that were repetitive or of low quality (identified by users with fewer than three interactions in their interaction history), leaving approximately 25K high-quality instructions. These instructions are mixed to create an instruction-tuning dataset to fine-tune the LLM.
5.1.4. Backbone Models
@@
  • Direct Recommendation. In the scenario of direct recommendation, We adopt four representative methods, including: MF (Koren et al., 2009), LightGCN (He et al., 2020), MixGCF (Huang et al., 2021), and SGL (Wu et al., 2021).
@@
  • Sequential Recommendation. Regarding sequential recommendation, we opt for three widely used models, including: SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), and CL4SRec (Xie et al., 2022).
@@
  • Rating Prediction. We consider the following classical models for rating prediction, including: DeepFM (Guo et al., 2017), NFM (He et al., 2017), DCN (Wang et al., 2017), AFM (Xiao et al., 2017), xDeepFM (Lian et al., 2018), and AutoInt (Song et al., 2019).
@@
We employ the LLaMA-2 7B version as the backbone LLM across all experiments, unless specifically mentioned otherwise. Our primary comparison is with the standard Instruction Fine-Tuning (IFT) method adopted in TALLRec (Bao et al., 2023) and InstructRec (Zhang et al., 2023c). For the rating prediction task, LLaMA-2 with IFT is used to directly predict the rating. For the top-k recommendation task, the tuned LLM is used to re-rank the list predicted by the backbone model, in accordance with (Hou et al., 2023), referred to as listwise ranking. Besides, we also adopt LLM for predicting the rating for each item and sort by the predicted scores, referred to as pointwise ranking.
5.1.5. Implementation Details.
@@
During training for LLaMA 2 (7B) with full-parameter tuning, we use a uniform learning rate of $2\times 10^{-5}$ and a context length of 2048, and we set the batch size as 16. Additionally, we use a cosine scheduler for three epochs in total with a 50-step warm-up period. To efficiently train the computationally intensive models, we simultaneously employ DeepSpeed training with ZeRO-3 stage (Rajbhandari et al., 2020) and flash attention (Dao et al., 2022). We trained the 7B model on 16 NVIDIA A800 80GB GPUs. For the inference stage, we employed the vLLM framework (Kwon et al., 2023) with greedy decoding, setting the temperature to 0. Only one GPU was utilized during the inference phase. We only evaluate the instruction-tuned LLaMA model for rating prediction task since it is not applicable for directly making top-$k$ recommendations.
[…]
@@
Table 4. Performance achieved by different methods in rating prediction task.
@@
5.2. Main Results (RQ1)
We conducted an extensive evaluation of our proposed Llama4Rec and the baseline methods on three datasets to assess the model’s performance under diverse recommendation scenarios. The experiment results for rating prediction, direct recommendation, and sequential recommendation are shown in Table 2, Table 3, and Table 4, respectively. We have the following key observations.
  • Llama4Rec consistently outperforms baseline methods in almost all scenarios, with particularly significant improvements observed in the direct recommendation task. Moreover, our findings reveal that direct instruction fine-tuning LLMs for recommendation tasks does not consistently yield promising performance. These results highlight the effectiveness of integrating LLMs into conventional recommendation models, underscoring the importance of incorporating the mechanism that utilizes instruction-tuned LLM to mutually augment and adaptively aggregate with conventional recommendation models.
  • In the scenario of the rating prediction task, while the instruction-tuned LLaMA model significantly underperforms when compared to conventional recommendation models, integrating the LLM yields a marked performance improvement. This suggests that the LLM and conventional recommendation models learn distinct aspects of information. Consequently, integrating the LLM with conventional recommendation models could enhance recommendation performance.
  • In the context of top-$k$ recommendations, Llama4Rec exhibits a more pronounced improvement for direct recommendations task. In addition, a more significant enhancement is observed on the Bookcrossing dataset, which can be attributed to the more fine-grained and distinguishable rating of the Bookcrossing dataset.
Table 5. Ablation study on key components of Llama4Rec on the ML-1M dataset.
@@ 5.2. Main Results (RQ1)
5.3. Ablation Study (RQ2)
We conducted an ablation study to analyze the contributions of different components in our model. Table 5 summarizes the results of the ablation studies across three variants on the ML-1M dataset. It is evident that the full model performs considerably better than all its variants, indicating that all the main components contribute significantly to overall performance improvement. Moreover, compared to the conventional model, the instruction-tuned LLM does not achieve superior results, underscoring the importance of model aggregation. We further analyze the specific impact of each component, and our observations are as follows:
@@ 5.3. Ablation Study (RQ2)
  • w/o Prompt Augmentation (w/o PA): In this variant, we remove the prompt augmentation component, a crucial element of the proposed framework. Experimental results demonstrate a significant degradation in model performance when this module is excluded, thereby validating its essential role. By employing the instruction-tuned LLM with prompt augmentation from prior knowledge by conventional recommendation models, we achieve an enhanced model performance, attributable to the capture of different aspects of information.
@@
5.4. Hyper-parameter Study (RQ3)
5.4.1. Analysis of Hyper-parameters $\alpha_{1}$ and $\alpha_{2}$
@@
Figure 3. Impact of hyper-parameters $\alpha_{1}$ and $\alpha_{2}$ on ML-1M dataset with backbone model LightGCN.
Figure 4. Performance comparison w.r.t different LLaMA-2 size for training Llama4Rec on the Bookcrossing dataset.
Figure 5. Performance comparison w.r.t different numbers of instructions for training Llama4Rec on the ML-1M dataset.
5.4.2. Analysis of Model Scaling.
@@
5.4.3. Analysis of Data Scaling.
@@
5.5. Further Discussion
In this part, we discuss about the computational efficiency and future improvements. In the Llama4Rec framework, additional training with augmented data is required, which may present a potential limitation. In the current experimental setup, we train a new model from scratch. However, this process could be optimized by continuing to train a previously tuned model, thereby reducing time costs. Additionally, in our experiment, we observed that training the LLaMA-2 7B model with around 25K instructions on 16 A800 GPUs with 2500 steps took approximately 1.94 hours. The inference time for each instruction averaged about 17 instructions per second, translating to a requirement of around 0.059 seconds per item for computation by a single A800 GPU.
@@ 5.5. Further Discussion
This training and inference duration significantly exceeds that of conventional recommendation models, highlighting the limitations of current LLM-based recommender systems. The substantial demand for computational resources also represents a significant challenge. Consequently, employing instruction LLMs for large-scale industrial recommender systems, such as those with millions of users, is presently impractical. However, future advancements in accelerated and parallel computing algorithms for language model inference could potentially reduce inference times and computation resources. This improvement might make the integration of LLMs into large-scale recommender systems feasible, especially by leveraging many GPUs for parallel computation.
6. Conclusion and future work

### v2

@@ IV Experiment
In this section, we present a thorough empirical evaluation to validate the effectiveness of our proposed framework. Specifically, our objective is to investigate whether the incorporation of our proposed Llama4Rec can enhance existing recommendation models. The overarching goal is to answer the following research questions:
@@
  • RQ3: How do different hyper-parameters impact the overall performance of the framework? RQ4: How does the instruction-tuned model compare to other LLMs, such as GPT? How would the model perform w.r.t. model and data scaling?
IV-A Experiment Setup
IV-A1 Dataset
Following [3], we rigorously evaluate the performance of our proposed framework by employing three heterogeneous, real-world datasets. MovieLens [^1] https://grouplens.org/datasets/movielens/ [13] serve as benchmark datasets in the realm of movie recommendation. We employ two variants of the dataset: MovieLens-100K (ML-100K) and MovieLens-1M (ML-1M). The former consists of approximately 100,000 user-item ratings, while the latter scales up to roughly 1,000,000 ratings. BookCrossing [^2] Due to the absence of timestamp data, we synthesize historical interactions through random sampling. [74] includes user-generated book ratings on a scale of 1 to 10, alongside metadata such as ‘Book-Author’ and ‘Book-Title’. We employ the LLM to augment the ‘director’ and ‘star’ features for ML-100K and ML-1M datasets, and augment the ‘genre’ and ‘page length’ features for the BookCrossing dataset. To ensure the data quality, we adopt the 5-core setting, where we filter unpopular users and items with fewer than five interactions for the BookCrossing dataset. The key characteristics of these datasets are delineated in the Table III.
TABLE III: Dataset Description.
@@ IV-A1 Dataset
IV-A2 Evaluation Metrics.
Aligning with [14, 49], for the top-$k$ recommendation task, we turn to two well-established metrics: Hit Ratio (HR) and Normalized Discounted Cumulative Gain (NDCG), denoted by H and N, respectively. In our experiments, $k$ is configured to be either 3 or 5 for a comprehensive evaluation, similar to the experiment setting in [66]. For instance, H@$k$ refers to the hit rate at position $k$, while N@$k$ denotes the normalized discounted cumulative gain (NDCG) at position $k$. In accordance with [9], we employ Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) as evaluation metrics for the rating prediction task.
TABLE IV: Performance achieved by different sequential recommendation methods. H@$k$ indicates the hit rate at position $k$, and N@$k$ represents normalized discounted cumulative gain (NDCG) at position $k$.
@@
IV-A3 Data Preprocessing.
Following prior works [66, 40], we adopt a leave-one-out evaluation strategy. More specifically, within each user’s interaction sequence, we choose the most recent item as the test instance. The item immediately preceding this serves as the validation instance, while all remaining interactions are used to constitute the training set. Moreover, regarding the instruction-tuning dataset construction, we randomly sampled 5K instructions for each of the three recommendation tasks (rating prediction, direct recommendation, and sequential recommendation) across three datasets (ML-100K, ML-1M, and BookCrossing), resulting in a total of 3 × 3 × 5K = 45K raw data. We eliminated instructions that were repetitive or of low quality (identified by users with fewer than three interactions in their interaction history), leaving approximately 25K high-quality instructions. These instructions are mixed to create an instruction-tuning dataset to fine-tune the LLM.
IV-A4 Backbone Models
@@
  • Direct Recommendation. In the scenario of direct recommendation, We adopt four representative methods, including: MF [22], LightGCN [14], MixGCF [19], and SGL [59].
@@
  • Sequential Recommendation. Regarding sequential recommendation, we opt for three widely used models, including: SASRec [20], BERT4Rec [49], and CL4SRec [63].
@@
  • Rating Prediction. We consider the following classical models for rating prediction, including: DeepFM [12], NFM [15], DCN [55], AFM [62], xDeepFM [26], and AutoInt [47].
@@
We employ the LLaMA-2 7B version as the backbone LLM across all experiments, unless specifically mentioned otherwise. Our primary comparison is with the standard Instruction Fine-Tuning (IFT) method adopted in TALLRec [3] and InstructRec [66]. For the rating prediction task, LLaMA-2 with IFT is used to directly predict the rating. For the top-k recommendation task, the tuned LLM is used to re-rank the list predicted by the backbone model, in accordance with [18], referred to as listwise ranking. Besides, we also adopt LLM for predicting the rating for each item and sort by the predicted scores, referred to as pointwise ranking.
IV-A5 Implementation Details.
@@
During training for LLaMA 2 (7B) with full-parameter tuning, we use a uniform learning rate of $2\times 10^{-5}$ and a context length of 2048, and we set the batch size as 16. Additionally, we use a cosine scheduler for three epochs in total with a 50-step warm-up period. To efficiently train the computationally intensive models, we simultaneously employ DeepSpeed training with ZeRO-3 stage [44] and flash attention [6]. We trained the 7B model on 16 NVIDIA A800 80GB GPUs. For the inference stage, we employed the vLLM framework [23] with greedy decoding, setting the temperature to 0. Only one GPU was utilized during the inference phase. We only evaluate the instruction-tuned LLaMA model for rating prediction task since it is not applicable for directly making top-$k$ recommendations.
We implement the models for rating prediction task using the DeepCTR-Torch [^3] https://github.com/shenweichen/DeepCTR-Torch library. For the top-$k$ recommendation task, we utilize the SELFRec [^4] https://github.com/Coder-Yu/SELFRec library [65] for implementation. As for the hyper-parameter settings, $\alpha_{1}$ and $\alpha_{2}$ are selected from $\{$0.1, 0.3, 0.5, 0.7, 0.9}\} respectively for all experiments. $\mathcal{C}$ is fixed to 1 and $k^{\prime}$ is set to 10. We repeat the experiment five times and calculate the average. We report the best results obtained when the ranking method is selected from pointwise and listwise ranking. For all experiments, the best results are highlighted in boldfaces. * indicates the statistical significance for $p\leq 0.05$ compared to the best baseline method based on the paired t-test. Improv. denotes the improvement of our method over the best baseline method.
IV-B Main Results (RQ1)
We conducted an extensive evaluation of our proposed Llama4Rec and the baseline methods on three datasets to assess the model’s performance under diverse recommendation scenarios. The experiment results for direct recommendation, sequential recommendation, and rating prediction are shown in Table II, Table IV, and Table V, respectively. We have the following key observations.
  • As shown in Table II, Llama4Rec consistently outperforms baseline methods in almost all scenarios, with particularly significant improvements observed in the direct recommendation task. Moreover, our findings reveal that direct instruction fine-tuning LLMs for recommendation tasks does not consistently yield promising performance. These results highlight the effectiveness of integrating LLMs into conventional recommendation models, underscoring the importance of incorporating the mechanism that utilizes instruction-tuned LLM to mutually augment and adaptively aggregate with conventional recommendation models.
  • We observe a performance decline when dealing with a larger number of candidates (i.e., $k$) in direct and sequential recommendation scenario, as the improvement for $k=5$ is usually less than for $k=3$. This suggests a limitation of the LLM, where ranking and recommending the correct items become more challenging as the candidate pool increases. This cause the performance drops in some cases. This may be attributed to the “lost-in-the-middle” problem [30], where LLMs tend to focus more on the beginning and ending of sequences while neglecting the middle portion. This behavior contributes to a decline in performance as the size of the candidate set increases. Besides, larger candidate sets inherently increase the complexity of the task, requiring the model to distribute its attention across more options. Smaller or less powerful LLMs may struggle to maintain this balance, leading to reduced performance. To address this issue, we propose several possible remedies. First, employing more powerful LLMs can help mitigate the problems associated with larger candidate sets, as these models are better equipped to handle longer sequences. Second, Splitting large candidate sets into smaller subsets (e.g., partitioning large $k$ into smaller chunks) and aggregating the results can help the model maintain focus and improve accuracy for higher $k$ values.
  • In the scenario of the rating prediction task in Table V, while the instruction-tuned LLaMA model significantly underperforms when compared to conventional recommendation models, integrating the LLM yields a marked performance improvement. This suggests that the LLM and conventional recommendation models learn distinct aspects of information. Consequently, integrating the LLM with conventional recommendation models could enhance recommendation performance.
  • In the context of direct recommendation and sequential recommendation recommendations in Table II and IV , Llama4Rec exhibits a more pronounced improvement for direct recommendations task. In addition, a more significant enhancement is observed on the Bookcrossing dataset, which can be attributed to the more fine-grained and distinguishable rating of the Bookcrossing dataset.
  • Besides, we discuss the computational efficiency. Additional training with augmented data is required in the Llama4Rec framework, which presents an extra cost. In the current experimental setup, we train a new model from scratch. However, this process could be optimized by continuing to train a previously tuned model, thereby reducing time costs. Additionally, in our experiment, we observed that training the LLaMA-2 7B model with around 25K instructions on 16 A800 GPUs with 2500 steps took approximately 1.94 hours. The inference time for each instruction averaged about 17 instructions per second, translating to a requirement of around 0.059 seconds per item for computation by a single NVIDIA A800 GPU.
@@ IV-B Main Results (RQ1)
TABLE V: Performance achieved by different methods in rating prediction task.
@@
TABLE VI: Ablation study on key components of Llama4Rec on the ML-1M and Bookcrossing dataset.
@@
IV-C Ablation Study (RQ2)
We conducted an ablation study to analyze the contributions of different components in our model. Table VI summarizes the results of the ablation studies across three variants on the ML-1M and Bookcrossing dataset. It is evident that the full model performs considerably better than all its variants, indicating that all the main components contribute significantly to overall performance improvement. Moreover, compared to the conventional model ( e.g., LightGCN ), the instruction-tuned LLM ( e.g., IFT ) does not achieve superior results, underscoring the importance of model aggregation. We further analyze the specific impact of each component, and our observations are as follows:
@@ IV-C Ablation Study (RQ2)
  • w/o Prompt Augmentation (w/o PA): In this variant, we remove the prompt augmentation component. Experimental results demonstrate a significant degradation in model performance when this module is excluded, thereby validating its essential role. By employing the instruction-tuned LLM with prompt augmentation from prior knowledge by conventional recommendation models, we achieve an enhanced model performance, attributable to the capture of different aspects of information.
@@
IV-D Hyper-parameter Study (RQ3)
@@
Figure 4: Impact of hyper-parameters $\alpha_{1}$ and $\alpha_{2}$ on ML-1M dataset with backbone model LightGCN.
IV-E Model Analysis (RQ4)
Figure 5: Performance comparison w.r.t different LLaMA-2 size for training Llama4Rec on the Bookcrossing dataset.
Figure 6: Performance comparison w.r.t different numbers of instructions for training Llama4Rec on the ML-1M dataset.
IV-E1 Analysis of Model Scaling
@@
IV-E2 Analysis of Data Scaling
@@
IV-E3 Comparison with the GPT Model
We compare our Llama4Rec with the GPT model, specifically, the GPT-3.5-turbo [^6] https://platform.openai.com/docs/models/gpt-3-5 model. We choose the performance metric hit ratio for comparison. A random selection of 100 listwise ranking task samples from the Bookcrossing dataset with backbone model CLSRec served as the benchmark for GPT model. The experimental result is shown in Figure 7. The experimental result shows Llama4Rec win over the GPT model in 45% of the tasks, tied in 25%, and fell short in 30% of the times. This impressive result emphasizes the crucial role of instruction tuning in aligning general-purpose LLMs specifically for recommendation tasks.
Figure 7: Comparison between our instruction-tuned Llama4Rec with the GPT-3.5-turbo model. Tie refers to cases where the proposed framework and GPT achieve identical performance metrics, specifically the hit ratio.
V Further Discussion
@@
Model Efficiency. This training and inference duration significantly exceeds that of conventional recommendation models, highlighting the limitations of current LLM-based recommender systems. The substantial demand for computational resources also represents a significant challenge. Consequently, employing instruction LLMs for large-scale industrial recommender systems, such as those with millions of users, is presently impractical. However, future advancements in accelerated and parallel computing algorithms for language model inference could potentially reduce inference times and computation resources. This improvement might make the integration of LLMs into large-scale recommender systems feasible, especially by leveraging many GPUs for parallel computation.
Linear Combination. In this work, we chose a heuristic-based linear interpolation method due to its simplicity and effectiveness, as evidenced by our experiments. Training a neural network to determine optimal aggregation weights would require substantial labeled data to accurately capture the relationship between input factors and performance, which is resource-intensive and impractical for this study. In future work, we plan to explore advanced techniques like genetic programming or reinforcement learning for better modeling of complex factor interactions and enhance nonlinear aggregation effects through dynamic weight optimization.
Hallucination. To ensure the accuracy of attributes generated by the LLM, we focused on well-known movies and books, utilizing their reliable metadata. For the evaluation process, we use a two-step approach. First, attributes are generated using the LLaMA model. Then, we employ another advanced LLM, DeepSeek-V3 [28], as an evaluator to assess the correctness of the generated attributes. This approach, referred to as the ”LLM-as-a-Judge” paradigm [73], allows us to systematically evaluate the alignment between generated attributes. In cases of disagreement between the evaluator and the generated attributes, we conduct manual evaluations to resolve conflicts, ensuring a higher degree of reliability and accuracy.
To maintain consistency in our analysis, we randomly sampled 50 items from each dataset and utilized greedy decoding during generation to reduce randomness. For ML-100K and ML-1M, we evaluated the generated ”Movie Director” attribute, while for BookCrossing, we evaluated the generated ”Book Genres” attribute. The quantitative results of this evaluation process are presented in Figure 8.
Figure 8: Performance of LLM-generated attributes compared to ground-truth data.
From Figure 8, we observe that the generated attributes demonstrate a high level of agreement with the ground-truth data, exceeding 80% across all datasets and reaching of 90% in ML-1M. However, two common types of errors are observed. First, in some cases, the Llama model honestly admits it does not know the answer. This behavior could serve as a useful indicator during model training, suggesting that the corresponding movie might not be widely known. Second, in other cases, the LLM hallucinates and provides entirely incorrect answers. While these errors are undesirable, they may serve as informative noise, potentially contributing to improved model robustness.
We attribute the high agreement observed to the extensive pretraining of the LLMs on vast amounts of world knowledge, which enables them to generate attributes with a high degree of correctness. However, smaller models are more prone to errors or hallucinations, while larger models generally exhibit better performance due to their increased parameter count, broader training data coverage, and reduced hallucination tendencies.
The world knowledge embedded in LLMs provides a strong foundation for attribute generation. However, to further improve accuracy and mitigate the risk of hallucination, we identify several promising future directions. These include integrating Retrieval-Augmented Generation (RAG) [25] to ground the model in external, verifiable knowledge sources, leveraging structured Knowledge Graphs [17], or employing next-generation, more powerful foundation models to mitigate these issues.
VI Conclusion and future work
@@
VII Acknowledgement
This work was supported in part by the Research Grants Council of the Hong Kong SAR under Grant GRF 11217823, 11216225, and Collaborative Research Fund C1042-23GF, the National Natural Science Foundation of China under Grant 62371411, InnoHK initiative, the Government of the HKSAR, Laboratory for AI-Powered Financial Technologies.